INAF queue observing, 2018 Mar 22-23

Observer: R. Carini, A. Rossi
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: B Rothberg on call
Instrument: MODS1+MODS2, LBC

Summary

At the sunset the sky is completed covered
instruments: all instruments work well Night duration (12\xB0 el twilight) 2:20am- 12:35pm UT = 10 h 15m

Observing time: h 40m (7%)

Weather time loss: 0h (88%)

technical time loss 30m *(5%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibrations arc MODS1+MODS2 red only     Caccianiga, Prandoni
Calibrations dark LUCI1 + LUCI2 see log     Antoniucci
Calibrations flat LUCI1 + LUCI2 see log     H Ks imaging flats Palazzi 31
Calibrations flat LUCI1 see log     HK,zJ Gilli Sarracco Antoniucci Mignoli, slit 0.75 and 1
telluric Antoniucci HD 24879 LUCI1 N1.8, zj, Hk   luci1.0057-0064  
Cal flats standard MODS1 dual grating, blue only, red only      
Cal flats standard MODS2 blue only      

Details

Calibrations MODS arc red only mods1.20180323.0001-0006; mods2r.20180323.0001-0006

LUCI dark luci1.20180323.0008-0012; luci2.20180323.0008-0012 dit= 10, ndit=1, normal, lir. Dr Tau
  • luci1.0013-0015; luci2.0013-0017 dit=25, ndit=1, normal, lir. Dr Tau
  • luci1.0018-0022; luci2.0018-0022 dit 4 ndit 3 integrated lir , Dr Tau
  • luci1.0023-0027; luci2.0023-0027 dit=30, ndit=1, normale, mer, XZ Tau
  • luci1.0028-0032; luci2.0028-0032 dit =10, ndit =1, normal, mer, XZ Tau
  • luci1.0033-0037; luci2.0033-0037 dit=14, ndit=1, integrated, lir, RW AUR
  • luci1.0038-0042; luci2.0038-0042 dit=7, ndit=1, integrated, lir, RW AUR
  • luci1.0043-0047; luci2.0043-0047 dit=2.75, ndit=1, nor, lir, RW AUR

02:24 the sky lightly better, we try to observe back-up targets. Slew to telluric for XZtau Antoniucci, HD 24879
  • LUCI2 error, no mask loaded, Dave Thomson will take a look
  • We will use LUCI1 only. created LUCI-1 only script, 10 min lost
  • 02:44 start acquisition
  • acq luci1.0053-0056
  • sci luci1.0057-0064

03:04 slew to the target XZ Tau (slit 1.0'')
  • 03:06 we cancel the preset, to fix the problem of LUCI2 the telescope has to be at the zenit

03:21 problem fixed, re start the science of XZ Tau
  • the target is too near to the moon, we have to change target.

03:30 slew to RW Aur (slit 0.75'')
  • cheking pointing
  • 03:36* start acquisition
  • acq luci1.0065-0067; luci2.0053-0055
  • 03:42 guide lost, too many clouds, we stop.

03:55 sky completed covered
  • 03:59 closed to clouds

calibrations arc MODS2 blue only mods2.0001-0006 problems with mods1 and mods2. We can not continue the calibrations!!!!!

LUCI
  • arc+flat luci1.0068-0083, luci2.0056-0072 HKspec, lambda=1.93, slit 1.0'', Gilli, Mignoli luci2 to be repeated
  • arc+flat luci1.0084-0099, ==luci2.0073-0087==HKspec, lambda=1.93, slit 0.75'', Saracco , to be repeated
  • arc+flat luci1.0100-0113, luci2.0088-00106 zjspec, lambda=1.175, slit 0.75'', Antoniucci, to be repeated
  • arc+flat luci1.0114-0127, luci2.0107-0120 zjspec, lambda=1.175, slit 1.0'', Antoniucci
  • ima flat luci1.0133-0152, luci2.0126-0145 H and Ks, Palazzi 31, GRB161007A

07:11 wind exceeds the limit of 21.6 m/s

07:53 start again the calibrations with MODS
  • flat for the standard in blue only with MODS2 ==mods2b.0009-0011== we take images but mods2 has problem in this channel, we can not continue in blue or dual grating.
  • Also the blue channel of MODS1 has problems, it not lock on, we can not do the calibrations dual grating
  • flat for the standard in red only mods1r.0007-0009 MODS1
  • also MODS2 red doesn't work!!!!

10:00 problems with mods1 fixed
  • flats for standard dual grating mods1r.0010-0012, mods1b.0001-0006

-- %USERSIG{RobertaCarini - 2018-03-23}%

Comments

 
Topic revision: r15 - 26 Mar 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback